OpenStudy (anonymous):

At the moment there are 54,210 tagged birds in a certain wildlife refuge. If exactly 20 percent of all birds in the refuge are tagged, what percent of the untagged birds must be tagged so that half of all birds in the refuge are tagged?

OpenStudy (anonymous):

nobody wants to help? wow

OpenStudy (anonymous):

i got this one :)

OpenStudy (anonymous):

it's 37 1/2% un-tagged birds are 80% half of all birds are 50% tagged birds are 20%

OpenStudy (anonymous):

(50-20)/80=30/80=3/8

OpenStudy (anonymous):

In fact, if we're going to use numbers, let's say there are 100 birds. This means that 20 are tagged and 80 are un-tagged. For 1/2 the birds to be tagged, we need 50 tagged birds. In other words, we need 30 additional tagged birds. So, from the 80 un-tagged, we need to tag 30 birds. 30/80 = 3/8 = 37.5% 9 mins. and no help thats crazy O.o